Complete parts (a) through (e) for the population data below. 1, 4, 7, 10 a. Find the mean, l, of the variable. u= (Type an iFind the sample mean x for each possible sample of size n=2. Х X 1,4 Sample Sample 4,7 1,7 4, 10 1, 10 7, 10 (Type integers oFind the sample mean x for each possible sample of size n= 3. х Sample 1,4,7 1, 4, 10 1, 7, 10 4, 7, 10 (Type integers or decFind the sample mean x for each possible sample of size n= 4. Sample Х 1, 4, 7, 10 (Type an integer or a decimal rounded to tc. Use the dotplots from part (b) to create one plot with the sampling distributions for each sample size and interpret yourd. For each of the possible sample sizes, find the probability that the sample mean will equal the population mean. n Probabi
Answer #1

a) µ= population mean=(1+4+7+10)/4=5.5          

b)

sample Xbar
1 1
4 4
7 7
10 10

c. 0 1 2 3 4 5 6 7 8 9 10 11

=============

sample Xbar Xbar
(1,4) 2.5 (4,7) 5.5
(1,7) 4 (4,10) 7
(1,10) 5.5 (7,10) 8.5

D. 0 1 2 3 4 5 6 7 8 9 10 11

============

sample Xbar
(1,4,7) 4
(1,4,10) 5
(1,7,10) 6
(4,7,10) 7

C. 0 1 2 3 4 5 6 7 8 9 10 11

==========

sample Xbar
(1,4,7,10) 5.5

OC. a a ארז 0 1 2 3 4 5 6 7 8 9 10 11

c)

OA. и na 1. n=2 n3 n=4 x 0 1 2 3 4 5 6 7 8 9 10 11

OC. As the sample size increases, the sampling distribution of the sample mean clusters closer to the population mean.

d)

n Probability
1 0
2 2/6 = 1/3
3 0
4 1

e)

n Probability
1 0
2 2/6 = 1/3
3 2/4=1/2